One angle of a triangle is 20 degrees. The second angle's measurement is four times the third. What is the measure of the third?
Answer by Edwin McCravy(20054)

One angle of a triangle is 20 degrees. The second angle's measurement is four times the third. What is the measure of the third?

Since we know it's a triangle with one angle, the first angle A, 
the second angle B, and third angle C, we know that all three angles 
of any triangle must have sum 180°, so

A + B + C = 180 

One angle
 (the first angle), A,
of a triangle is 20 degrees.

So

A = 20

The second angle's
, B's,
measurement is four times
 C,
the third".

 So

B = 4C

So we can take the first equation

A + B + C = 180

and substitute 20 for A and 4C for B

20 + 4C + C = 180

    20 + 5C = 180

         5C = 180 - 20

         5C = 160

          C = 32°

That is the measure of the third,
but if you had been asked for B it
would have been calculated from

B = 4C
B = 4(32°) = 128°

Then you could check the answer by checking
the original words using:

First angle = A = 20°,
Second angle = B = 128°
Third angle = 32°

>>...One angle of a triangle is 20 degrees...<<

That checks because the first angle is 20°

>>...The second angle's measurement is four times the third...<<

That checks because 128° is 4x32°
Edwin